OpenStudy (anonymous):

If √x = 8 , then x =?? show work

OpenStudy (anonymous):

\[x = 64\]

OpenStudy (anonymous):

square both sides :O

OpenStudy (anonymous):

Or, try a slightly cooler approach using logarithms: \[x^{1/2} = 8\]\[x^{1/2} = 2^{3}\]\[\log_{2}(x^{1/2}) = 3\]\[(1/2)\log_{2}x = 3\]\[\log_{2}x = 6\]\[x = 2^{6}\]\[x = 6\]

OpenStudy (anonymous):

*Last line should read "x = 64"